# A Survey on Mixture of Experts

### v1

@@ 4.4. Training & Inference Scheme
To mitigate the training overhead associated with vision MoE transformer models, the Residual Mixture-of-Experts (RMoE) approach (Wu et al., 2022) commences with training a dense, non-MoE model on the upstream task, followed by an efficient fine-tuning stage to transition into a MoE model. This process reveals that directly inheriting expert weights from a pre-trained non-MoE model’s FFN can lead to suboptimal performance, necessitating an alignment between the MoE and FFN outputs during the fine-tuning phase. Similarly, Dua et al. (Dua et al., 2022) advocate for initially training a dense model, subsequently introducing sparsity by incorporating a randomly initialized gating module, and further training the model’s feed-forward layers under sparsity conditions—specifically, by updating the weights locally within each compute node rather than averaging gradients across nodes.
@@
Nie et al. (Nie et al., 2021) present EvoMoE, an efficient end-to-end MoE training framework. EvoMoE decouples the joint learning of experts and the sparse gate, emphasizing the acquisition of foundational knowledge through a single expert at the inception of training. Subsequently, it spawns multiple diverse experts and advances the diversification of experts by training with the novel Dense-to-Sparse gate (DTS-Gate). The DTS-Gate initially operates as a dense activation of all experts, then progressively and adaptively constricting to route tokens to a reduced number of experts. A similar strategy is employed in the development of the MoE-LLaVA (Lin et al., 2024) large vision-language model, which commences with a dense model, subsequently multiplies the feedforward network (FFN) to create expert initializations, and proceeds to train exclusively the MoE layers, while keeping the remaining model components static.
@@
Komatsuzaki et al. (Komatsuzaki et al., 2022) highlight the efficiency of sparse models in terms of quality and computational cost, yet acknowledge their significant data requirements and the expense of training from scratch at scale. To address this, they introduce a scheme termed ”sparse upcycling,” which leverages pre-existing training investments by initializing a sparsely activated MoE model from a pre-trained dense checkpoint. This involves transferring all parameters—and optionally their associated optimizer states—from the original checkpoint, with the exception of the MoE gating network parameters, which are not present in the dense model. Notably, the new MoE layers are populated with identical copies of the original dense model’s FFN layers, and the gating mechanism weights are initialized randomly. A critical obstacle in model upcycling is the initial performance decrease resulting from structural modifications to a trained network. To mitigate this performance regression during upcycling, the researchers proposes normalizing the gate combine weights for each token to 1. This approach is grounded in the notion that, in the dense model, each token was processed by a singular ”expert” FFN. While this normalization proved beneficial for upcycled vision models, it was found to be detrimental to the performance of upcycled language models.
@@
Building upon the sparse upcycling technique (Komatsuzaki et al., 2022), the Skywork-MoE model (Tianwen Wei, 2024) leverages the foundational architecture of its pre-developed Skywork-13B model (Wei et al., 2023), adopting its dense checkpoints as a foundation for initial states. Their empirical evidence indicates that the decision between sparse upcycling and training from scratch should be informed by both the performance of available dense checkpoints and the MoE-specific training resources, as models trained from scratch consistently surpass their upcycled counterparts in performance. The study observes a decline in average expert similarity throughout the training of upcycled MoEs, suggesting a diversification of experts emerges during the process. Importantly, the Skywork-MoE analysis reveals that models with greater expert similarity tend to underperform, establishing expert similarity as a potential diagnostic tool during MoE training for upcycled models. Conversely, the expert similarity in models trained from scratch remains minimal, implying that non-uniform expert initialization promotes diversification.
@@
Pan et al. (Pan et al., 2024) posit that the parameter inefficiency observed in MoE models stems from conventional sparse training methodologies, where only a select group of experts is engaged and refined for each input token. To counteract this, they introduce a hybrid framework for MoE models, denoted as DS-MoE, which integrates dense training (activating all experts) with sparse inference (sparse expert activation) to achieve higher computation and parameter efficiency. Notably, DS-MoE maintains activation for all self-attention experts (MoA (Zhang et al., 2022b)) during inference but selectively activates FFN experts, reflecting the observation that self-attention layers manifest considerably less sparsity compared to FFN layers.
@@
Chen et al. (Chen et al., 2022d) introduce SMoE-Dropout, an innovative plug-and-play training framework, which initially modularizes the FFN into a sequence of smaller FFNs then employs a random policy parameterized by fixed weights to route token to k experts with the largest response. Progressively, the framework activates an increasing number of experts, preventing overfitting to the amounts of used network capacity during training. MoEfication (Zhang et al., 2022a) investigates various strategies for expert construction in T5 models, including random splitting, parameter clustering, and building co-activation graphs. MoEBERT (Zuo et al., 2022) implements an importance-based method for adapting FFNs into experts within BERT models. LLaMA-MoE (Team, 2023) conducts an extensive examination of different expert construction methodologies, ultimately proposing a straightforward random division approach that partitions the parameters of FFNs into non-overlapping experts.
4.4.2. Sparse-to-Dense.
@@
Switch Transformer (Fedus et al., 2022b) studies the distillation of large sparse models into smaller dense counterparts to achieve parameter efficiency for deployment. The study reports that initializing the corresponding weights of dense model from non-expert layers of MoE model modestly enhances performance, facilitated by the consistent dimension of non-expert layers. Furthermore, an improvement in distillation is observed when employing a mixture of 0.25 for the teacher probabilities and 0.75 for the ground truth label. Leveraging both methods, the distillation preserves approximately 30% of the sparse model’s quality gains using only about 5% of the parameters. Similarly, Xue et al. (Xue et al., 2022a) address the challenges of overfitting, deployment difficulty, and hardware constraints associated with sparse MoE models. Drawing inspiration from human learning paradigms, they propose a new concept referred to as ’knowledge integration’ aimed at creating a dense student model (OneS) that encapsulates the expertise of a sparse MoE model. Their framework first implements knowledge gathering, explored through a variety of methods such as summation, averaging, top- Knowledge Gathering, and their Singular Value Decomposition Knowledge Gathering. Then, they refine the dense student model by knowledge distillation to mitigate noise introduced from the knowledge gathering. The OneS model retains 61.7% of the MoE’s benefits on ImageNet and 88.2% on NLP datasets. Further investigations into MoE model distillation are also conducted by other researchers (Rajbhandari et al., 2022; Costa-jussà et al., 2022).
@@
Chen et al. (Chen et al., 2022b) highlight the challenges associated with deploying MoE models on resource-constrained platforms, such as cloud or mobile environments. Observing that only a fraction of experts contribute significantly to MoE fine-tuning and inference, they propose a method for the progressive elimination of non-essential experts. This approach retains the advantages of MoE while simplifying the model into a single-expert dense structure for the target downstream task. Similarly, ModuleFormer (Shen et al., 2023c) applies a comparable pruning technique, removing task-unrelated experts for a lightweight deployment. Huang et al. (Huang et al., 2023) separate the training and inference stages for Vision Transformers (ViTs). They substitute certain FFNs in the ViT with custom-designed, efficient MoEs during training. These MoEs assign tokens to experts using a random uniform partition and incorporate Experts Weights Averaging (EWA) on these MoEs at the end of each iteration. After training, the MoEs are converted back to FFNs through averaging of expert weights, thus reverting the model to its original dense ViT for inference.
4.4.3. Expert Models Merging.
@@
Li et al. (Li et al., 2022) introduce the Branch-Train-Merge (BTM) algorithm, a method for the communication-efficient training of language models (LMs). BTM independently trains a set of expert LMs (ELMs), each tailored to a specific domain within the training corpus, such as scientific or legal text. These ELMs, which operate without shared parameters, can be ensembled or parameter-averaged at inference to coalesce into a singular LM. Expanding on this concept, Sukhbaatar et al. (Sukhbaatar et al., 2024) present Branch-Train-MiX (BTX), designed to combine the strengths of BTM and Mixture-of-Experts while addressing their respective limitations. BTX maintains separate training for multiple expert LLMs, akin to BTM, but subsequently integrates these experts within a unified MoE model. Specifically, it consolidates the FFNs from all ELMs into a singular MoE module at each layer, with a gating network determining the appropriate FFN expert for each token. Other components, such as the self-attention layers from ELMs, are merged by averaging their weights. The resulting model then undergoes MoE-finetuning on all the combined data to enable the gate to effectively mix the FFN experts.
@@
Wang et al. (Wang et al., 2023) point that while the emergence of Foundation Models made it easier to obtain expert models tailored to specific tasks, the heterogeneity of data at test time necessitates more than a single expert. Accordingly, they explore the Fusion of Experts (FoE) challenge, which aims to integrate outputs from expert models that provide diverse but complementary insights into the data distribution, formulating it as an instance of supervised learning.
4.5. Derivatives
@@
Building upon the principles of algorithm design highlighted earlier, numerous studies have drawn inspiration from the Mixture of Experts (MoE) framework, proposing a range of MoE variants. We categorize these innovative models as derivatives of the MoE. For instance, Xue et al. (Xue et al., 2022b) introduced WideNet, an approach that increases model width by substituting the feed-forward network (FFN) with an MoE layer while maintaining shared trainable parameters across Transformer layers, except for the normalization layers. Subsequently, Tan et al. (Tan et al., 2023) presented the Sparse Universal Transformer (SUT), an efficient enhancement of the Universal Transformer, which is characterized by parameter-sharing across its layers. SUT incorporates a Sparse Mixture of Experts and a novel stick-breaking-based dynamic halting mechanism, thus reducing computational complexity without compromising parameter efficiency or generalization capabilities. Moreover, the traditional MoE models often employ discrete matching between experts and tokens (Shazeer et al., 2017; Lepikhin et al., 2020; Du et al., 2022; Artetxe et al., 2021; Zhou et al., 2022a; Fedus et al., 2022b; Zoph et al., 2022), a practice associated with training instability and uneven expert utilization. Addressing these challenges, Antoniak et al. (Antoniak et al., 2023) innovatively proposes the Mixture of Tokens (MoT), which blends tokens from different examples before presenting them to the experts. Thus, MoT enables the model to benefit from a wider array of token-expert combinations.
@@
Recently, the MoE’s principle of assigning specialized knowledge to individual experts has been adapted to parameter-efficient fine-tuning (PEFT). Choi et al. (Choi et al., 2023) propose the sparse mixture-of-prompts (SMoP), a method that utilizes a gating mechanism to train multiple short soft prompts, each adept at processing distinct subsets of data. This addresses the inefficiencies encountered with long soft prompts during prompt tuning. The MoE framework has also been integrated into lifelong learning (LLL), which seeks to facilitate continuous learning from an ongoing stream of data. The Lifelong-MoE model (Chen et al., 2023c) dynamically expands model capacity by adding experts with regularized pretraining, effectively mitigating the issue of catastrophic forgetting (Kirkpatrick et al., 2017) typically associated with straightforward fine-tuning. In a recent development, the MoE concept of conditional computation has been further refined to optimize resource allocation in transformer-based language models (LMs). The Mixture-of-Depths (MoD) (Raposo et al., 2024) employs a binary gating network to decide whether a token should be processed by a given Transformer layer. As a result, MoD transformers can dynamically allocate computational resources (FLOPs) to specific sequence positions, achieving a lower overall FLOP footprint compared to vanilla or MoE-based transformers.
@@
In summary, the evolution of MoE derivatives reveals a trend where models either integrate the conditional computation aspect of the gating mechanism or merge the MoE structure with various tasks achieved by assigning specialized knowledge to individual experts, such as aforementioned prompt tuning (Choi et al., 2023) and lifelong learning (Chen et al., 2023c) with MoE, demonstrating the versatility and adaptability of the MoE architecture across different domains.
5. System Design of Mixture of Experts
While Mixture of Experts (MoE) has been increasingly leveraged to enhance the capabilities of large language models, its adoption introduces new challenges to existing training and inference systems, due to the inherently sparse and dynamic nature of its computational workload. Since the introduction of expert parallelism by GShard(Lepikhin et al., 2020), which implements parallel gating and expert computation by dispatching partitioned local tokens with the load balancing limit of expert capacity, this paradigm has emerged as a fundamental strategy to facilitate the efficient scaling of MoE models. This approach can be viewed as an augmentation of data parallelism (Rajbhandari et al., 2020; Ren et al., 2021; Rajbhandari et al., 2021), where each expert in an MoE layer is assigned to a distinct device, while all non-expert layers are duplicated across devices. As depicted in Figure 8(a), the process flow of expert parallelism consists of the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. In general, the input size for GEMM needs to be large enough to achieve optimal utilization and throughput of computing device necessitates. Therefore, input encode is employed to aggregate the input tokens of a same expert into a contiguous memory space, as determined by the token-expert mapping from gate routing. Subsequently, the All-to-All dispatch is employed to send the input tokens to their corresponding experts across the distributed devices. Following the localized computation by the experts, the inverse process - All-to-All combine and output decode, reinstates the original data layout according to the the gating indices.
Furthermore, the synergy of expert parallelism (Fedus et al., 2022b; Hwang et al., 2023a; Singh et al., 2023; Ma et al., 2022; Zheng et al., 2022) with other existing parallel strategies (tensor (Shoeybi et al., 2019; Smith et al., 2022; Narayanan et al., 2021), pipeline (Huang et al., 2019; Narayanan et al., 2019; Qi et al., 2023), sequence parallelism (Li et al., 2021; Korthikanti et al., 2023; Jacobs et al., 2023)) has been investigated to enhance the scalability and efficiency of MoE models within large-scale distributed environments. As shown in Figure 8, we illustrate several examples pf hybrid parallelism, encompassing (b) data + expert + tensor parallelism (Fedus et al., 2022b; Rajbhandari et al., 2022; Singh et al., 2023; Hwang et al., 2023a; Zhai et al., 2023), (c) data + expert + pipeline parallelism (He et al., 2022; Zhai et al., 2023; Hwang et al., 2023a), (d) expert + tensor parallelism (Tianwen Wei, 2024). It is imperative to recognize that the choice of distributed parallelism strategies influences a complex interplay between computation efficiency, communication overhead, memory occupation, potentially affected by various hardware configurations. Consequently, the deployment strategies for practical applications necessitate nuanced trade-offs and bespoke designs tailored to specific use-case scenarios.
In the subsequent discussion, we delineate the challenges introduced by MoE models from computation, communication, and storage aspects, concurrently reviewing existing research addressing these issues. Table 4 shows an overview of the open-source MoE frameworks.
@@ 5. System Design of Mixture of Experts
| OpenMoE (Xue et al., 2024) | Colossal-AI | ✓ | ✓ | https://github.com/hpcaitech/ColossalAI | 38K | |
| ScatterMoE (Tan et al., 2024) | Mila Quebec | ✓ | https://github.com/shawntan/scattermoe | 140 | ||
| Megablocks (Gale et al., 2023) | Stanford University | ✓ | https://github.com/stanford-futuredata/megablocks | 1.1K | ||
| Tutel (Hwang et al., 2023a) | Microsoft | ✓ | ✓ | https://github.com/microsoft/tutel | 672 | |
| SE-MOE (Shen et al., 2022) | Baidu | ✓ | ✓ | ✓ | https://github.com/PaddlePaddle/Paddle | 21K |
| HetuMoE (Nie et al., 2022) | Peking University | ✓ | ✓ | https://github.com/PKU-DAIR/Hetu | 236 | |
| Deepspeed-MoE (Rajbhandari et al., 2022) | Microsoft | ✓ | ✓ | https://github.com/microsoft/DeepSpeed | 33K | |
| FastMoE (He et al., 2021) | Tsinghua University | ✓ | ✓ | https://github.com/laekov/fastmoe | 1.4K | |
| Fairseq (Ott et al., 2019; Artetxe et al., 2021) | Meta | https://github.com/facebookresearch/fairseq/tree/moe | 29K | |||
| Mesh-TensorFlow (Shazeer et al., 2018) | https://github.com/tensorflow/mesh | 1.6K | ||||
5.1. Computation

### v2

@@ 4.4.1 Dense-to-Sparse
To mitigate the training overhead associated with vision MoE transformer models, the Residual Mixture-of-Experts (RMoE) approach [122] commences with training a dense, non-MoE model on the upstream task, followed by an efficient fine-tuning stage to transition into a MoE model. This process reveals that directly inheriting expert weights from a pretrained non-MoE model’s FFN can lead to suboptimal performance, necessitating an alignment between the MoE and FFN outputs during the fine-tuning phase. Similarly, Dua et al. [123] advocate for initially training a dense model, subsequently introducing sparsity by incorporating a randomly initialized gating module, and further training the model’s feed-forward layers under sparsity conditions—specifically, by updating the weights locally within each compute node rather than averaging gradients across nodes.
@@
Nie et al. [62] present EvoMoE, an efficient end-to-end MoE training framework. EvoMoE decouples the joint learning of experts and the sparse gate, emphasizing the acquisition of foundational knowledge through a single expert at the inception of training. Subsequently, it spawns multiple diverse experts and advances the diversification of experts by training with the novel Dense-to-Sparse gate (DTS-Gate). The DTS-Gate initially operates as a dense activation of all experts, then progressively and adaptively constricting to route tokens to a reduced number of experts. A similar strategy is employed in the development of the MoE-LLaVA [106] large vision-language model, which commences with a dense model, subsequently multiplies the feed-forward network (FFN) to create expert initializations, and proceeds to train exclusively the MoE layers, while keeping the remaining model components static.
@@
Komatsuzaki et al. [49] highlight the efficiency of sparse models in terms of quality and computational cost, yet acknowledge their significant data requirements and the expense of training from scratch at scale. To address this, they introduce a scheme termed ”sparse upcycling,” which leverages pre-existing training investments by initializing a sparsely activated MoE model from a pretrained dense checkpoint. This involves transferring all parameters—and optionally their associated optimizer states—from the original checkpoint, with the exception of the MoE gating network parameters, which are not present in the dense model. Notably, the new MoE layers are populated with identical copies of the original dense model’s FFN layers, and the gating mechanism weights are initialized randomly. A critical obstacle in model upcycling is the initial performance decrease resulting from structural modifications to a trained network. To mitigate this performance regression during upcycling, the researchers propose normalizing the gating scores for each input token to 1, which are used to combine the outputs of multiple experts. This approach is grounded in the notion that, in the dense model, each token was processed by a singular “expert” FFN. While this normalization proved beneficial for upcycled vision models, it was found to be detrimental to the performance of upcycled language models. In the first introduction of sparse MoE by [27], the softmax function was applied to the selected top- gating values, which normalizes the combination gating scores to 1, formulated as . However, subsequent LLMs [28, 36, 38] with MoE have evolved to apply the softmax function across all potential gating values before isolating the top- subset, formulated as .
@@
Building upon the sparse upcycling technique [49], the Skywork-MoE model [67] leverages the foundational architecture of its pre-developed Skywork-13B model [193], adopting its dense checkpoints as a foundation for initial states. Their empirical evidence indicates that the decision between sparse upcycling and training from scratch should be informed by both the performance of available dense checkpoints and the MoE-specific training resources, as models trained from scratch consistently surpass their upcycled counterparts in performance. The study observes a decline in average expert similarity throughout the training of upcycled MoEs, suggesting a diversification of experts emerges during the process. Importantly, the Skywork-MoE analysis reveals that models with greater expert similarity tend to underperform, establishing expert similarity as a potential diagnostic tool during MoE training for upcycled models. Conversely, the expert similarity in models trained from scratch remains minimal, implying that non-uniform expert initialization promotes diversification.
@@
Pan et al. [64] posit that the parameter inefficiency observed in MoE models stems from conventional sparse training methodologies, where only a selected group of experts is engaged and refined for each input token. To counteract this, they introduce a hybrid framework for MoE models, denoted as DS-MoE, which integrates dense training (activating all the experts) with sparse inference (sparse expert activation) to achieve higher computation and parameter efficiency. Notably, DS-MoE maintains activation for all self-attention experts (MoA [82]) during inference but selectively activates FFN experts, reflecting the observation that self-attention layers manifest considerably less sparsity compared to FFN layers.
@@
Chen et al. [124] introduce SMoE-Dropout, an innovative plug-and-play training framework, which initially modularizes the FFN into a sequence of smaller FFNs then employs a random policy parameterized by fixed weights to route token to k experts with the largest response. Progressively, the framework activates an increasing number of experts, preventing overfitting to the amounts of used network capacity during training. MoEfication [50] investigates various strategies for expert construction in T5 models, including random splitting, parameter clustering, and building co-activation graphs. MoEBERT [125] implements an importance-based method for adapting FFNs into experts within BERT models. LLaMA-MoE [51] conducts an extensive examination of different expert construction methods, ultimately proposing a straightforward random division approach that partitions the parameters of FFNs into non-overlapping experts. Emergent MoEs (EMoE) [126] splits certain FFN layers of the original model into MoE layers with clustering-based experts construction and avg-k gating, which ameliorates the parameter updating during fine-tuning and can even be abandoned afterward to preserve the original model architecture.
4.4.2 Sparse-to-Dense
@@
Switch Transformer [36] studies the distillation of large sparse models into smaller dense counterparts to achieve parameter efficiency for deployment. The study reports that initializing the corresponding weights of dense model from non-expert layers of MoE model modestly enhances performance, facilitated by the consistent dimension of non-expert layers. Furthermore, an improvement in distillation is observed when employing a mixture of 0.25 for the teacher probabilities and 0.75 for the ground truth label. Leveraging both methods, the distillation preserves approximately 30% of the sparse model’s quality gains using only about 5% of the parameters. Similarly, Xue et al. [52] address the challenges of overfitting, deployment difficulty, and hardware constraints associated with sparse MoE models. Drawing inspiration from human learning paradigms, they propose a new concept referred to as ‘knowledge integration’ aimed at creating a dense student model (OneS) that encapsulates the expertise of a sparse MoE model. Their framework first implements knowledge gathering, explored through a variety of methods such as summation, averaging, top- Knowledge Gathering, and their Singular Value Decomposition Knowledge Gathering. Then, they refine the dense student model by knowledge distillation to mitigate noise introduced by the knowledge gathering. The OneS model retains 61.7% of the MoE’s benefits on ImageNet and 88.2% on NLP datasets. Further investigations into MoE model distillation are also conducted by other researchers [66, 78].
@@
Chen et al. [53] highlight the challenges associated with deploying MoE models on resource-constrained platforms, such as cloud or mobile environments. Observing that only a fraction of experts contribute significantly to MoE fine-tuning and inference, they propose a method for the progressive elimination of non-essential experts. This approach retains the advantages of MoE while simplifying the model into a single-expert dense structure for the target downstream task. Similarly, ModuleFormer [73] applies a comparable pruning technique, removing task-unrelated experts for a lightweight deployment. Huang et al. [127] separate the training and inference stages for Vision Transformers (ViTs). They substitute certain FFNs in the ViT with custom-designed, efficient MoEs during training. These MoEs assign tokens to experts using a random uniform partition and incorporate Experts Weights Averaging (EWA) on these MoEs at the end of each iteration. After training, the MoEs are converted back to FFNs through averaging of expert weights, thus reverting the model to its original dense ViT for inference. He et al. [128] propose a unified framework for MoE model compression, encompassing two strategies: Expert Slimming, which compresses individual experts via pruning and quantization, and Expert Trimming, which structurally removes unimportant experts.
4.4.3 Expert Models Merging
@@
Li et al. [129] introduce the Branch-Train-Merge (BTM) algorithm, a method for the communication-efficient training of language models (LMs). BTM independently trains a set of expert LMs (ELMs), each tailored to a specific domain within the training corpus, such as scientific or legal text. These ELMs, which operate without shared parameters, can be ensembled or parameter-averaged at inference to coalesce into a singular LM. Expanding on this concept, Sukhbaatar et al. [54] present Branch-Train-MiX (BTX), designed to combine the strengths of BTM and Mixture-of-Experts while addressing their respective limitations. BTX maintains separate training for multiple expert LLMs, akin to BTM, but subsequently integrates these experts within a unified MoE model. Specifically, it consolidates the FFNs from all ELMs into a singular MoE module at each layer, with a gating network determining the appropriate FFN expert for each token. Other components, such as the self-attention layers from ELMs, are merged by averaging their weights. The resulting model then undergoes MoE fine-tuning on all the combined data to enable the gate to effectively mix the FFN experts.
@@
Wang et al. [130] point out that while the emergence of Foundation Models made it easier to obtain expert models tailored to specific tasks, the heterogeneity of data at test time necessitates more than a single expert. Accordingly, they explore the Fusion of Experts (FoE) challenge, which aims to integrate outputs from expert models that provide diverse but complementary insights into the data distribution, formulating it as an instance of supervised learning.
4.5 Derivatives
@@
Building upon the principles of algorithm design highlighted earlier, numerous studies have drawn inspiration from the Mixture of Experts (MoE) framework, proposing a range of MoE variants. We categorize these innovative models as derivatives of the MoE. For instance, Xue et al. [58] introduced WideNet, an approach that increases model width by substituting the feed-forward network (FFN) with an MoE layer while maintaining shared trainable parameters across Transformer layers, except for the normalization layers. Subsequently, Tan et al. [59] presented the Sparse Universal Transformer (SUT), an efficient enhancement of the Universal Transformer, which is characterized by parameter-sharing across its layers. SUT incorporates a Sparse Mixture of Experts and a novel stick-breaking-based dynamic halting mechanism, thus reducing computational complexity without compromising parameter efficiency or generalization capabilities. Moreover, the traditional MoE models often employ discrete matching between experts and tokens [27, 28, 35, 103, 94, 36, 37], a practice associated with training instability and uneven expert utilization. To address these challenges, Antoniak et al. [56] innovatively proposes the Mixture of Tokens (MoT), which blends tokens from different examples before presenting them to the experts. Thus, MoT enables the model to benefit from a wider array of token-expert combinations.
@@
Recently, the MoE’s principle of assigning specialized knowledge to individual experts has been adapted to parameter-efficient fine-tuning (PEFT). Choi et al. [60] propose the sparse mixture-of-prompts (SMoP), a method that utilizes a gating mechanism to train multiple short soft prompts, each adept at processing distinct subsets of data. This addresses the inefficiencies encountered with long soft prompts during prompt tuning. The MoE framework has also been integrated into lifelong learning (LLL), which seeks to facilitate continuous learning from an ongoing stream of data. The Lifelong-MoE model [55] dynamically expands model capacity by adding experts with regularized pretraining, effectively mitigating the issue of catastrophic forgetting [194] typically associated with straightforward fine-tuning. In a recent development, the MoE concept of conditional computation has been further refined to optimize resource allocation in transformer-based language models (LMs). The Mixture-of-Depths (MoD) [57] employs a binary gating network to decide whether a token should be processed by a given Transformer layer. As a result, MoD transformers can dynamically allocate computational resources (FLOPs) to specific sequence positions, achieving a lower overall FLOP footprint compared to vanilla or MoE-based transformers.
@@
In summary, the evolution of MoE derivatives reveals a trend where models either integrate the conditional computation aspect of the gating mechanism or merge the MoE structure with various tasks achieved by assigning specialized knowledge to individual experts, such as aforementioned prompt tuning [60] and lifelong learning [55] with MoE, demonstrating the versatility and adaptability of the MoE architecture across different domains.
@@
OpenMoE [38] | Colossal-AI | ✓ | ✓ | https://github.com/hpcaitech/ColossalAI | 38K | |
ScatterMoE [140] | Mila Quebec | ✓ | https://github.com/shawntan/scattermoe | 140 | ||
Megablocks [139] | Stanford University | ✓ | https://github.com/stanford-futuredata/megablocks | 1.1K | ||
Tutel [132] | Microsoft | ✓ | ✓ | https://github.com/microsoft/tutel | 672 | |
SE-MoE [133] | Baidu | ✓ | ✓ | ✓ | https://github.com/PaddlePaddle/Paddle | 21K |
HetuMoE [136] | Peking University | ✓ | ✓ | https://github.com/PKU-DAIR/Hetu | 236 | |
Deepspeed-MoE [66] | Microsoft | ✓ | ✓ | https://github.com/microsoft/DeepSpeed | 33K | |
FastMoE [131] | Tsinghua University | ✓ | ✓ | https://github.com/laekov/fastmoe | 1.4K | |
Fairseq [195, 103] | Meta | https://github.com/facebookresearch/fairseq/tree/moe | 29K | |||
Mesh-TensorFlow [196] | https://github.com/tensorflow/mesh | 1.6K |
5 System Design of Mixture of Experts
While Mixture of Experts (MoE) has been increasingly leveraged to enhance the capabilities of large language models, its adoption introduces new challenges to existing training and inference systems, due to the inherently sparse and dynamic nature of its computational workload. GShard [28] introduces expert parallelism that implements parallel gating and expert computation by dispatching partitioned local tokens with load balancing limit of expert capacity. Since then, expert parallelism has emerged as a fundamental strategy to facilitate efficient scaling of MoE models. This approach can be viewed as an augmentation of data parallelism [197, 198, 199], where each expert in an MoE layer is assigned to a distinct device, while all non-expert layers are duplicated across devices. As depicted in Figure 8(a), the process flow of expert parallelism consists of the following sequential operations: gate routing, input encode, All-to-All dispatch, expert computation, All-to-All combine, and output decode. In general, the input size for general matrix multiply (GEMM) needs to be large enough to achieve optimal utilization and throughput that computing device necessitates. Therefore, input encode is employed to aggregate the input tokens of a same expert into a contiguous memory space, as determined by the token-expert mapping from gate routing. Subsequently, the All-to-All dispatch is employed to send the input tokens to their corresponding experts across the distributed devices. Following the localized computation by the experts, the inverse process—All-to-All combine and output decode—reinstates the original data layout according to the gating indices.
Furthermore, the synergy of expert parallelism [36, 132, 135, 200, 201] with other existing parallel strategies (tensor [202, 203, 204], pipeline [205, 206, 207], sequence parallelism [208, 209, 210]) has been investigated to enhance the scalability and efficiency of MoE models in large-scale distributed environments. As shown in Figure 8, we illustrate several examples of hybrid parallelism, encompassing (b) data + expert + tensor parallelism [36, 66, 135, 132, 138], (c) data + expert + pipeline parallelism [134, 138, 132], and (d) expert + tensor parallelism [67]. It is imperative to recognize that the choice of distributed parallelism strategies influences a complex interplay between computation efficiency, communication overhead, memory occupation, potentially affected by various hardware configurations. Consequently, the deployment strategies for practical applications necessitate nuanced trade-offs and bespoke designs tailored to specific use-case scenarios.
In the subsequent discussion, we delineate the challenges introduced by MoE models from computation, communication, and storage aspects, concurrently reviewing existing research addressing these issues. Table IV shows an overview of the open-source MoE frameworks.
5.1 Computation